I don't understand how to solve this linear equation -2b+6=-10
Answer by Theo(13342)

-2b + 6 = -10
multiply both sides of the equation by -1 to get:
2b - 6 = 10
add 6 to both sides of the equation to get:
2b = 10 + 6
simplify to get:
2b = 16
divide both sides of the equation by 2 to get:
b = 8
that should be your answer after you confirm that it is good.
substitute 8 for b in the original equation of:
-2b + 6 = -10 to get:
-2(8) + 6 = -10
simplify to get:
-16 + 6 = -10
combine like terms to get:
-10 = -10
this is a true statement so the value of 8 for b is good.
your answer is:
b = 8